| ID |
Date |
Author |
Subject |
Text |
 |
|
248
|
Tue, 3.06.2025 22:31:28 |
T.Czopowicz | Preparations for 'neutrino run' 2025 | 1. Vdrift measurements stopped ~7am due to
an error in SY4527 crate (CA claims that
write operation failed, when it actually |
|
|
247
|
Sun, 1.06.2025 15:54:16 |
T.Czopowicz | Preparations for 'neutrino run' 2025 | Added TPC temperatures to grafana.
Added cooling water temperatures and pressures
to grafana. |
|
|
246
|
Sat, 31.05.2025 19:36:42 |
T.Czopowicz | Preparations for 'neutrino run' 2025 | Fixed issue with status=0 in GAS.
Added temperatures, etc. to grafana |
|
|
245
|
Fri, 30.05.2025 10:46:11 |
T.Czopowicz | Final fix for GAS | Added additional field "raw" for ADC values |
|
|
244
|
Thu, 29.05.2025 10:43:55 |
T.Czopowicz | Preparations for 'neutrino run' 2025 | - removed obsolete AiRecord and switched
to tmux (teslameter, betaerg, DIP, GAS, PTB200,
S7-200, etherbox, Kikusui, zentro)
|
|
|
243
|
Thu, 19.10.2023 18:41:39 |
T.Czopowicz | New developments | 1. sound alarm in Control Room
2. mRPC ToF HVin grafana properly displayed |
|
|
Draft
|
Sun, 9.07.2023 11:13:04 |
T.Czopowicz | CAMAC ADC failure | The CAMAC module, that measures temperatures
from the drift velocity monitors failed yesterday
~8am.
|
6x |
|
241
|
Tue, 30.10.2018 15:46:52 |
T.Czopowicz | camac ADC readout moved to epics | CAMAC ADCs (temperatures used by old NA9 system)
are read out by epics |
|
|
240
|
Sun, 8.07.2018 12:43:38 |
T.Czopowicz | na61dcs1 resterted | Due to repeating problem:
gas application halts with dialog saying
something about not enough memory
|
|
|
239
|
Wed, 4.07.2018 21:10:04 |
T.Czopowicz | EPICS support for ptb2000 barometers installed | Pressure in counting house rack 24 is now
read by epics (both barometers) |
|
|
238
|
Wed, 4.07.2018 19:45:40 |
T.Czopowicz | New etherbox readout | Etherbox in now read by EPICS application
Edit (6.07.2018 13:14:17):
|
|
|
237
|
Wed, 4.07.2018 15:59:45 |
T.Czopowicz | Bug in LV fixed | VME board sending commands to heat cooling
water had bad address.
Fixed and tested. |
|
|
236
|
Tue, 3.07.2018 21:34:37 |
T.Czopowicz | New readout of temperatures in MTPC and VTPC | New, EPICS version od the readout installed
in ~/DCS/S7-200 |
|
|
235
|
Tue, 26.06.2018 11:03:01 |
T.Czopowicz | Renaming channels in EPICS | Changed
FTPC2a:[..] => FTPC2:[...]
|
|
|
234
|
Fri, 9.03.2018 15:55:05 |
T.Czopowicz | PLCs and pulsers + dhcp | From now on, PLCs (measuting temperates in
MTPCs and VTPCs) and pulsers obtain IP via
bootstrap protocol from standard DHCP server |
|
|
233
|
Mon, 11.12.2017 10:00:18 |
B. Maksiak | DCS DB mirror at WUT down | Mon, 11.12.2017 10:01:28
DCS database went down this night because
of disk filesystem going to read-only mode. |
|
|
232
|
Wed, 31.05.2017 10:46:14 |
T. Czopowicz | LV check | VME controllers' firmware upgraded and connected
to the new PCI card inside the new na61dcs
computer. |
|
|
231
|
Mon, 12.12.2016 15:46:43 |
B. Maksiak | PCI Caenet card dismounted from na61dcs1 | The Caenet PCI card has been dismounted from
na61dcs1 computer. It will be mounted in
the new na61dcs machine. |
|
|
230
|
Sat, 3.12.2016 08:42:47 |
B. Maksiak | Troubleshooting | In case of run crashing with message: Jumper
settngs (0x*) in * miniheader is incorrect
for CBx, MBy.
|
|
|
229
|
Thu, 24.11.2016 15:00:37 |
T.Czopowicz | New switch in MTPC hut | A small switch installed inside MTPC hut |
|
|
228
|
Mon, 21.11.2016 09:25:54 |
T.Czopowicz | Betaerg readout | If a flow meter is switched off, it shows
0 flow. |
|
|
227
|
Mon, 21.11.2016 09:24:32 |
T.Czopowicz | Modification to raspberryPi software | All EPICS applications on raspberryPis are
now under one IOC.
Also they start with the system via init |
|
|
226
|
Mon, 21.11.2016 09:14:37 |
T.Czopowicz | LV VME boards address check | VME connector<->address correspondence has
been checked for MTPCs.
|
|
|
225
|
Wed, 26.10.2016 00:48:44 |
H.-J. Mathes | Switched to stable software release | The software (links) for the TPC gas system
were switched (at 2016/10/25 14:30) to the
current stable/production version in /home/na61dcs/src/GasMonitor/Scripts/Misc/run_EPICS.py.
|
|
|
224
|
Sun, 2.10.2016 20:27:31 |
T.Czopowicz | Small fixes since the last noghts' failures | VTPC-1 trigger level changed from -300 to
-170 mV.
|
|
|
Draft
|
Sun, 2.10.2016 20:27:00 |
T.Czopowicz | | |
|
|
222
|
Thu, 22.09.2016 02:20:20 |
B. Maksiak | High voltage communication stopped | Communication with high voltage crate sy527
stopped. Java application that steers high
voltages was not responding.
|
|
|
221
|
Mon, 25.07.2016 18:49:02 |
A.Aduszkiewicz | Gas application hanged | The gas application hanged with an error concerning
PREMA and GPIB (see attached screenshot).
After stopping and starting again it seems |
|
|
220
|
Wed, 20.07.2016 21:53:57 |
T.Czopowicz | trigger2db bug fix | Due to small coding mistake trigger2db couldn't
reconnect to trigger server after rebooting.
|
|
|
219
|
Thu, 3.03.2016 00:09:36 |
T.Czopowicz, B.Maksiak | na61dcs1 system reinstalled | System on the na61dcs1 computer has been reinstalled
(Debian 8) |
|
|
218
|
Sun, 28.02.2016 00:18:28 |
B. Maksiak, T. Czopowicz | Birth of na61dcs3 | The na61dcs3 machine has been created from
recycled parts:
[LIST]
|
|
|
217
|
Fri, 26.02.2016 16:44:54 |
T.Czopowicz, B.Maksiak | na61dcst system reinstalled | na61dcst computer system reinstalled (Debian
8). |
|
|
216
|
Fri, 26.02.2016 16:43:41 |
T.Czopowicz, B.Maksiak | CAEN PCI card driver updated for newest kernel version | CAEN PCI card driver updated for newest kernel
version.
That was the only obstacle to reinstalling |
|
|
215
|
Fri, 26.02.2016 16:42:34 |
T. Czopowicz, B. Maksiak | Tidying DCS switch | Network cables connected to DCS switch were
revised. Two cables were removed because
they were not connected to any device on |
|
|
214
|
Thu, 25.02.2016 10:55:39 |
T.Czopowicz, B.Maksiak | PSD position monitoring fixed | Monitoring of the PSD position has been improved.
Instead of supplying Raspberry Pi with 5V
via the GPIO pins,
|
|
|
213
|
Thu, 25.02.2016 10:08:32 |
T. Czopowicz, B. Maksiak | DCS workspace moved | The DCS workspace has been moved. This included:
[LIST]
[*] Moving DCS computers to the desk with |
|
|
212
|
Mon, 30.11.2015 02:34:20 |
B. Maksiak | Drift velocity measurement in VTPC2 dropped | VDrift measurement dropped from 1.4 to 1.1
cm/us.
|
|
|
211
|
Fri, 27.11.2015 12:46:56 |
T.Czopowicz | Trigger -> DB | Changes to trigger2db:
- S1=0 is sent to the DB (requested by Szymon),
- receiving reply from the DB implemented,
|
|
|
210
|
Sun, 22.11.2015 10:56:40 |
B. Maksiak | Raspberry Pis installation | Last week new raspberry pis were installed
to measure pressure, temperature and drift
velocity in GTPC and future FTPCs. They are |
|
|
209
|
Fri, 20.11.2015 17:56:41 |
B.Maksiak | Relative Humidity measurement in VTPC hut is back | Relative Humidity measurement in VTPC hut
is fixed. It was a problem of connnection. |
|
|
208
|
Thu, 19.11.2015 16:34:31 |
T.Czopowicz | Magnetic measurement stopped | Magnetic field measurement for both magnets
stopped as we do not plan to use magnets
during this run. |
|
|
207
|
Thu, 19.11.2015 16:33:11 |
T.Czopowicz | Pressure sensors | 4 Raspberry Pi machines ready to measure pressure
in drift monitors. |
|
|
206
|
Wed, 18.11.2015 16:53:55 |
T.Czopowicz | VD measurement | Looking for peak in averaged "close" result
begins from x=150 to prevent from taking
into account the noise at x=100. |
|
|
205
|
Sat, 14.11.2015 18:49:18 |
A. Aduszkiewicz | Labview hanged | Drift velocity measurement in VTPC-1 hanged
around 13:50. Labview works very slowly.
Restarted the gas application.
|
|
|
204
|
Fri, 16.10.2015 17:08:18 |
O. Wyszynski, B. Maksiak | LV card changed | 16.10.2015 17:08:37 (BM) Low Voltage card
202 was broken. It was changed for a new
one. |
|
|
203
|
Wed, 9.09.2015 11:45:44 |
Wojciech Dominik | E/p calibration of the MTPCs | MTPC-R E/p calibration measurement started
MTPC-L E/p cal meas done yesterday at noon,
but should be repeated because of MTPC hut |
|
|
200
|
Sat, 21.03.2015 13:53:50 |
T. Czopowicz | Magnetic field -> DAQ | Magnetic field measurements results are now
sent to DAQ (and to event header) |
|
|
199
|
Tue, 3.03.2015 12:59:40 |
B. Maksiak | DCS network changes | Preparation to install new teslameters.
A small network switch has been installed |
|
|
198
|
Tue, 20.01.2015 11:11:39 |
A. Aduszkiewicz | Humidity measurement in the VTPC hut | Humidity measurement in the VTPC hut was dead
since the end of the last run (Dec 15, 2014).
The temperature measurement was still working, |
|
|
197
|
Mon, 8.12.2014 14:48:51 |
B. Maksiak | DCS LabVIEW application moved | The Gas DCS LabVIEW application was moved
from na61dcst to na61dcs1 computer. Mainly
because there is no longer need from shifters |
|
|
196
|
Fri, 24.10.2014 15:26:46 |
T.Czopowicz | Picoscope connected to LabVIEW | It is now possible to measure drift velocity
with picoscope instead of LeCroy oscilloscope. |
 |
|
195
|
Thu, 23.10.2014 14:12:53 |
T.Czopowicz | PSD counting house connected to the DCS switch | PS Counting House connected to the DCS switch.
Now computers can connect to the DCS independently
from the CERN network. |
|
|
194
|
Sun, 19.10.2014 09:53:34 |
T.Czopowicz | na61dcs2 - system reinstalled | System reinstalled on na61dcs2.
Newer kernel was needed to install picoscope
- new oscilloscope to measure drift velocity. |
|
|
193
|
Sat, 20.09.2014 23:41:56 |
B. Maksiak | Java monitors lost connection | The java monitor stopped to taking measurements
in two periods. First period (20.09, 12-20;
see bottom-right plot) is caused by small |
|
|
192
|
Sat, 20.09.2014 23:31:20 |
B. Maksiak | PSD position monitor installed and connected | PSD position monitor has been connected. The
device stays at PSD hut. Connected via ethernet
cable directly to DCS switch. Cable lays |
|
|
191
|
Thu, 11.09.2014 08:56:53 |
T.Czopowicz | VTPC LV | Cable for VTPC VL power supplies mounting
finished.
|
|
|
190
|
Wed, 10.09.2014 12:16:07 |
T.Czopowicz | HV connection tested | Connection with HV module tested.
HV on one MTPC-L sector set to 10
V to test communication. |
|
|
189
|
Tue, 9.09.2014 15:56:23 |
A.Aduszkiewicz, T.Czopowicz | drift velocity DCS | The VTPC-1 and MTPC-L drift velocity stopped
to work yesterday. Today we found out one
of the scopes, Lecroy 6841, is broken. With |
|
|
188
|
Sun, 17.08.2014 16:50:39 |
A.Aduszkiewicz | Temperature sensors restared | The boxes reading the temperature sensors
in the VTPC hut and the MTPC hut were restarted.
This seems to fix a problem with missing |
|
|
187
|
Thu, 24.01.2013 23:48:07 |
A.Aduszkiewicz | E/p scan | E/p scan completed in all 4 TPCs.
Scan started at 8AM. First interrupted by |
|
|
186
|
Sat, 19.01.2013 13:24:35 |
A.Aduszkiewicz | measurements hanged | na61dcs2 computer hanged in the night for
several hours. It was fixed by Krzysztof,
but meanwhile several measurements went wrongly, |
|
|
185
|
Wed, 16.01.2013 10:35:56 |
A.Aduszkiewicz | pressure measurement hanged | The pressure measurement hanged around 9AM.
Fixed within half hour by powering the barometer
off and on. |
|
|
184
|
Fri, 11.01.2013 11:12:37 |
H.-J. Mathes | GAS Monitor readout | Already yesterday evening around 20:00, started
full readout, write every 30 seconds to na61db.
Before (since shutdown in December 2012) |
|
|
183
|
Mon, 7.01.2013 18:16:06 |
A.Aduszkiewicz | pressure measurement hanged | I realised that the pressure measurement hanged.
Attempted to fix it by doing the power cycle. |
|
|
182
|
Fri, 14.12.2012 13:05:38 |
A.Aduszkiewicz | VTPC-2/MTPC-R gas air conditioning broke | Today around 3AM air conditioning in the gas
rack VTPC-2/MTPC-R broke. It cools down to
16 C, although set to 20. It may cause some |
|
|
181
|
Wed, 12.12.2012 14:26:47 |
A.Aduszkiewicz | E/p scan done | E/p scan done successfully in all 4 TPCs. |
|
|
180
|
Sat, 8.12.2012 16:21:56 |
A.Aduszkiewicz | MTPC-R compressor failure | The compressor failure in MTPC-R occured.
Around 11:00 it was replaced and the gas
circulation was restarted around 13:00. This |
|
|
179
|
Tue, 4.12.2012 12:27:02 |
B. Maksiak | New temperature sensor (PSD cal.) in Measurement Monitor | 03.12.2012 15:45
PSD calorimeter temperature sensor has been
installed and included to DCS Java Measurement |
|
|
178
|
Mon, 3.12.2012 10:59:15 |
B. Maksiak | DCS crash | DCS crashed. Drift velocity readouts was "nan" |
|
|
177
|
Sun, 2.12.2012 15:44:25 |
A.Aduszkiewicz | E/p scan | E/p scan started in all 4 chambers.
03.12.2012 10:40 Scan finished successfully
in all 4 TPCs. Updated vdrift parameter. |
|
|
176
|
Sat, 1.12.2012 15:45:05 |
A.Aduszkiewicz | vd "TPC on" | As the drift voltage is full, I changed the
Labview DCS setting to "TPC on", so the monitor
voltage will follow the TPC voltage. |
|
|
175
|
Thu, 29.11.2012 11:41:05 |
A.Aduszkiewicz | E/p scan | 2012.11.29 11:40:42: started in all 4 TPCs
13:00 Labview crashed. Restarting the measurement
2012.11.30 11.30 Scan finished successfully. |
|
|
174
|
Fri, 23.11.2012 17:28:48 |
A.Aduszkiewicz | PTB-200 pressure hanged | The pressure measurement hanged on Nov 7.
The pressure displayed on the unit was correct,
but apparently the communication with PC |
|
|
173
|
Wed, 3.10.2012 08:17:42 |
A.Aduszkiewicz | E/p scan | 2012/10/03 8:17 started in 4 TPCs
2012/10/03 13:48 successful measurement in
VTPC-2 and both MTPCs (updated slope parameter). |
   |
|
172
|
Fri, 21.09.2012 01:41:26 |
MKir | E/p scan done | DCST machine was killed (fan pressing on a
switch in the power distributor in the control
house) during the evening shift, DCS failed |
   |
|
171
|
Thu, 20.09.2012 14:31:33 |
MKir | E/P scan | The chambers have been up to full drift voltage
for ~3h.
E/P scan started ~14:30.
|
|
|
170
|
Mon, 6.08.2012 15:37:53 |
A.Aduszkiewicz | Gas DCS | 1. Set the "TPC on or off" to 'off' as a default
value
2. Switched off 3 Gas DCS crates |
|
|
169
|
Sun, 5.08.2012 19:12:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
168
|
Sat, 4.08.2012 12:59:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
167
|
Sat, 4.08.2012 10:34:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
166
|
Wed, 1.08.2012 16:27:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
165
|
Tue, 31.07.2012 03:34:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
164
|
Mon, 30.07.2012 00:26:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
163
|
Sun, 29.07.2012 15:26:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
162
|
Sun, 29.07.2012 00:07:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
161
|
Sat, 28.07.2012 16:39:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
160
|
Sat, 28.07.2012 11:27:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
158
|
Fri, 27.07.2012 02:34:34 |
B. Maksiak | Changes in Gas Application | Changed default state of button "TPC ON or
OFF?" to ON. |
|
|
157
|
Thu, 26.07.2012 06:54:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
156
|
Wed, 25.07.2012 04:52:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
155
|
Wed, 25.07.2012 00:18:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
154
|
Tue, 24.07.2012 10:16:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
153
|
Mon, 23.07.2012 17:56:09 |
A.Aduszkiewicz | VTPC-1 salve T sensor moved | VTPC-1 salve temperature sensor attached to
the magnet in order to monitor the magnet
temperature during warming up. It should |
|
|
152
|
Mon, 23.07.2012 01:25:21 |
A.Aduszkiewicz | E/p scan | 2012.07.23 01:25 E/p scan started
2012.07.23 12:42 Updated vdrift parameter
in VTPCs, but the application crashed before |
|
|
151
|
Sat, 21.07.2012 14:58:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
150
|
Fri, 20.07.2012 17:50:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
149
|
Fri, 20.07.2012 11:48:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
148
|
Fri, 20.07.2012 02:52:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
147
|
Wed, 18.07.2012 18:25:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|
|
146
|
Wed, 18.07.2012 17:59:01 |
Gas DCS | Gas System Crashed | System crashed. Restarted |
|